OpenStudy (anonymous):

Let f(x) = x2 + 3x - 4 and g(x) = x + 5. Find f(x) · g(x).

OpenStudy (anonymous):

just multiply (x^2+3x-4)*(x+5)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

x^3+5x^2+3x^2+15x-4x-20 x^3+8x^2+11x-20

OpenStudy (anonymous):

\[x^{3}+8x^{2}+11x-20?\]

OpenStudy (austinl):

\(f(x)=x^2+3x-4\\ g(x)=x+5\) \(f(x)\times g(x)=(x^2+3x-4)(x+5)=\) \(=x^3+5x+3x^2+15x-4x-20\\ =\color{blue}{x^3+3x^2+16x-20}\)

OpenStudy (austinl):

You have to distribute each of the terms individually.

OpenStudy (anonymous):

I'm so sorry, my internet cut out and I wasn't able to reply

OpenStudy (austinl):

No worries, might not be your internet though. This site has been known to wig out on occasion.